How much water must be evaporated from 240 gallons of a 3% salt solution to produce a 5% salt solution?
Answer by mananth(16949)

0.03*240=0.05(240-x)
7.2=12-0.05x
-4.8 =-0.05x
-4.8/-0.05=x
x=96 gallons